I'm in fourth grade and need help with this problem. Is 1/4 of 3, 3/4, 3 divided by 4, 4 divided by 3 the same and why? THANK YOU!
Answer by RAY100(1637)

"of" means multiply,,,,(1/4) of 3 ,,,,means (1/4) *3 = 3/4
.
think of this as 3 sets of 1/4 or 1/4 +1/4 +1/4 = 3/4
.
.
OR 3/4,,,,,3* 1/4 =3/1*1/4= (3*1)/(1*4) = 3/4
.
think of this as dividing 3 into 4 equal pieces,,,,first divide each of
.
the three circles,,into 4 parts,,,,this will give 12 total parts,,each (1/4)
.
now,,divide the 12 pieces into 4 groups,,,,each group would have 3 pieces (each 1/4)
.
or 4 sets of 3/4.
.
.
The reciprocal comes into play as we divide
.
Starting with 3*1/4 = 3/1 * 1/4 = 3/4
.
to divide,,,"invert and multiply" 2nd term
.
3/(1/4) = (3/1) / (1/4) = (3/1)*(4/1) = 12 pieces,,,,,(as we did above)
.
.
To better understand this,, draw three circles,,,divide each one
.
into 4 pieces....each piece is (1/4) and there are 12 total.
.
.
Review this several times,,until you understand
.
Thanks for your hard work
.